Find the Laplace transform, if it exists, of each of the following functions

(a) f(t) = eat

(b) f(t) = 1

(c) f(t) = t 


1
Expert's answer
2021-09-23T16:50:16-0400

(a)


F(s)=L(f(t))=L(eat)=0eatestdtF(s)=L(f(t))=L(e^{at})=\displaystyle\int_{0}^{\infin}e^{at}e^{-st}dt

=1salimA[e(as)t]A0=1sa(01)=-\dfrac{1}{s-a}\lim\limits_{A\to\infin}[e^{(a-s)t}]\begin{matrix} A \\ 0 \end{matrix}=-\dfrac{1}{s-a}(0-1)




=1sa=\dfrac{1}{s-a}

(b)


F(s)=L(f(t))=L(1)=0(1)estdtF(s)=L(f(t))=L(1)=\displaystyle\int_{0}^{\infin}(1)e^{-st}dt




=1slimA[est]A0=1s(01)=-\dfrac{1}{s}\lim\limits_{A\to\infin}[e^{-st}]\begin{matrix} A \\ 0 \end{matrix}=-\dfrac{1}{s}(0-1)

=1s=\dfrac{1}{s}

(c)


F(s)=L(f(t))=L(t)=0(t)estdtF(s)=L(f(t))=L(t)=\displaystyle\int_{0}^{\infin}(t)e^{-st}dt

=1slimA[test+1sest]A0=1s(00+01s)=-\dfrac{1}{s}\lim\limits_{A\to\infin}[te^{-st}+\dfrac{1}{s}e^{-st}]\begin{matrix} A \\ 0 \end{matrix}=-\dfrac{1}{s}(0-0+0-\dfrac{1}{s})

=1s2=\dfrac{1}{s^2}